Can you please help me solve: Solve using the addition method.
3/7 x + 5/9 y = 27
1/9 x + 2/7 y = 7
Thank you!

Answer by ptaylor(2198)


3/7 x + 5/9 y = 27-------------eq1
1/9 x + 2/7 y = 7 ------------eq2
Thank you!
Multiply each term in eq2 by 27/7 and we get:
(where did I get 27/7?? I solved a little equation (1/9)*x=3/7; x=27/7)
(27/7)(1/9)x+(27/7)(2/7)=7*(27/7) simplify
3/7 x +54/49 y=27------------eq2a

subtract eq2a from eq1 (or add (-eq2a) to eq1):
(5/9-54/49)y=27-27=0
y=0
substitute y=0 into eq1
3/7 x+(5/9)*0=27
3/7x=27
x=(7*27)/3=63
substitute x=63, y=0 into eq2
(1/9)*63+0=7
7=7
substitute y=0,x=63 into eq1
(3/7)*63+0=27
27=27

Hope this helps---ptaylor